### v2

# Generalized Bayesian Inference for Causal Effects using Covariate Balancing Procedure

Abstract
In observational studies, the propensity score plays a central role in estimating causal effects of interest. The inverse probability weighting (IPW) estimator is commonly used for this purpose. However, if the propensity score model is misspecified, the IPW estimator may produce biased estimates of causal effects. Previous studies have proposed some robust propensity score estimation procedures. However, these methods require considering parameters that dominate the uncertainty of sampling and treatment allocation. This study proposes a novel Bayesian estimating procedure that necessitates probabilistically deciding the parameter, rather than deterministically. Since the IPW estimator and propensity score estimator can be derived as solutions to certain loss functions, the general Bayesian paradigm, which does not require the considering the full likelihood, can be applied. Therefore, our proposed method only requires the same level of assumptions as ordinary causal inference contexts. The proposed Bayesian method demonstrates equal or superior results compared to some previous methods in simulation experimentss, and is also applied to real data, namely the Whitehall dataset.
Keywords: covariate balancing, general Bayes, inverse probability weighting, M-estimator, propensity score
1 Introduction
Adjusting for covariates (or confounders) is important in estimating the causal effects in observational studies. In the case where all covariates requiring adjustment are observed, it is possible to obtain a consistent estimator for causal effects, commonly known as “no unmeasured confounding[9].” In this context, the propensity score plays a central role in estimating causal effects[25]. However, because the propensity score is usually unknown, it must be estimated using appropriate procedures, and estimators may be biased if the propensity score model is misspecified[18]. Therefore, constructing a valid model is crutial for accurate estimation.
Imai and Ratkovic[16] propose the method known as the “Covariate Balancing Propensity Score” (CBPS), which does not involve modeling the propensity score directly, but, instead estimates the propensity score satisfying a moment restriction. Specifically, the moment condition comprises covariates that are desired to have a balanced distribution between the treatment groups, and the propensity score model is estimated to achieve covariate balance. The CBPS estimation heavily relies on selecting the appropriate moment restriction. Imai and Ratkovic recommend estimating CBPS based on second-order moments of covariates to achieve balance. Huang et al.[11] argue that achieving balance for higher-order moments may be necessary to obtain more accurate estimates of causal effects. More recently, to achieve a more precise covariate balance, several methods have been proposed. For example, kernel balancing methods[34, 8] aim to improve the balance by using kernel functions. Additionally, a method that satisfies the balancing of the characteristic function of the covariates has been proposed[12]. However, their methods require exact fulfillment of covariate balancing conditions, implying an equational constraint on a moment condition is necessary. Zubizarreta[36] argues that achieving an exact balance is challenging because the balance condition depends on the law of large numbers, which is applicable only in infinite samples. In other words, the condition does not account for the uncertainty derived from random sampling and treatment allocation. To address this issue, Zubizarreta proposes the “stable balancing weights” (SBW) method[36], which ensures a covariate balance up to a specified upper bound $\lambda$ that can be determined by researchers. In other words, the covariate balancing condition is subject to the subjective decision of the researcher. It is worth noting that Tan[31] demonstrates that the propensity score obtained using a logistic regression model that satisfies the same condition as the SBW method can be considered as the solution for a loss function. This result is highly relevant and critical for our proposed procedure.
The SBW-based method depends on moment conditions for covariate balance and the upper bound parameter $\lambda$, which determines the acceptable range of deviation for the covariate balancing condition. As mentioned in the previous paragraph, the upper bound is subjectively determined by the researcher. This study adopts the idea of Bayesian methods to propose a methodology that relaxes the decision process of the parameter. Specifically, we leverage the conclusion that a causal effect estimator (we consider the inverse probability weighting [IPW] estimator) and the SBW-based method can be described as solutions for a loss function. Then, we proceed to specify prior distributions for each model. Additionally, to achieve the mentioned objective, we specify a prior distribution for $\lambda$. To derive the posterior distribution, which plays a central role in Bayesian inference, we employ the general Bayes procedure [3]. Using the Bayesian inference procedure, our proposed method can construct credible intervals and perform statistical tests within the Bayesian framework, not only for the causal effects of interest but also for the parameter $\lambda$.
To the best of our knowledge, this study is the first to estimate the IPW estimator from a Bayesian perspective with a covariate balancing condition. The remainder of this paper is organized as follows. In Section 2, we introduces the notations and significant previous works. We specifically demonstrate that the SBW-based propensity score estimation method and IPW estimator can be explained through certain loss functions. In Section 3, we propose a Bayesian-based estimation procedure using the General Bayes paradigm. Section 4 presents simulation experiments to evaluate the performance of our proposed method against that of some competitors. In Section 5, we apply the methods to real data: Whitehall dataset.
2 Preliminaries
In this section, we introduce covariate balancing methods and their properties, specifically CBPS[16] and the SBW method[36, 31]. Furthermore, we demonstrate that the IPW estimator, used for estimating causal effects, can be considered as the solution for a loss function. Although this information is well-known, it forms a crucial foundation for the subsequent discussions in this section.
[…]
In observational studies, the propensity score is usually unknown[17]. Therefore, we need to specify the model and derive the predictor of the propensity score for the subsequent analysis. Generally, the propensity score model is assumed to be the logistic regression model. However, model misspecification can lead to crucial bias in the main analysis, that is, when estimating the causal effects. To overcome the problem, Imai and Ratkovic propose CBPS, which involves estimating the parameter $\bm{\alpha}$ of the propensity score model $e(\bm{X}_{i};\bm{\alpha})$ by satisfying the following moment restriction[16]:
$\frac{1}{n}\sum_{i=1}^{n}\left(\frac{A_{i}}{e(\bm{X}_{i};\bm{\alpha})}-\frac{1-A_{i}}{1-e(\bm{X}_{i};\bm{\alpha})}\right)g(\bm{X}_{i})=\bm{0},$
where $g(\cdot)$ is a measurable finite dimensional ($L$ dimension) function related to the covariates; for instance,
$g(\bm{X}_{i})=\left(X_{i1},\dots,X_{ip},X_{i1}^{2},\dots,X_{ip}^{2}\right).$
In the context of missing data, Zubizarreta[36] proposes the SBW method. To adapt it to the causal inference context, the SBW-based method can be considered as estimating $\bm{\alpha}$ using the following moment restriction (cf. Tan[31]):
$\frac{1}{n}\left|\sum_{i=1}^{n}\left(\frac{A_{i}}{e(\bm{X}_{i};\bm{\alpha})}-\frac{1-A_{i}}{1-e(\bm{X}_{i};\bm{\alpha})}\right)g_{j}(\bm{X}_{i})\right|<\lambda\ \ \ (j=1,2,\dots,L),$ (2.1)
where $\lambda$ is a hyper-parameter that determines the upper bound of the covariate balance. Zubizarreta demonstrates that the upper bound $\lambda$ influences the bias of the mean of interest for the marginal expectation of outcome [36]. Theoretically, Wang and Zubizarreta show that this upper bound must converge to 00 with a certain probability order to ensure accuracy [32]. Furthermore, they discuss a bootstrap sample-based algorithm for determining $\lambda$.
2.1 Loss Function for the Propensity Score Estimation using Covariate Balancing Procedure
Tan[31] derives an important conclusion, stating that the SBW-based method can be considered as the solution of a loss function when the propensity score model is assumed to be the logistic regression model. Specifically, the loss function is described as follows:
$\ell_{A}(\bm{\alpha})=\sum_{i=1}^{n}\left[A_{i}\exp\left\{-\bm{\alpha}^{\top}g(\bm{X}_{i})\right\}+(1-A_{i})\bm{\alpha}^{\top}g(\bm{X}_{i})+(1-A_{i})\exp\left\{\bm{\alpha}^{\top}g(\bm{X}_{i})\right\}-A_{i}\bm{\alpha}^{\top}g(\bm{X}_{i})\right]$ (2.2)
with the following penalty term:
$\lambda\|\bm{\alpha}\|_{1}:=\lambda\sum_{j=1}^{L}|\alpha_{j}|.$ (2.3)
Note that the relationship between (2.1), and (2.2) and (2.3) is derived from the duality of a constrained optimization problem.
2.2 Loss Function for the IPW Estimator
After estimating the propensity score, for instance through the loss function (2.2), we proceed to estimate the interesting causal effects using methods such as the IPW estimator, which is described as follows:
$\hat{\tau}=\hat{\theta}_{1}-\hat{\theta}_{0}=\frac{\sum_{i=1}^{n}\frac{A_{i}Y_{i}}{e_{i}(\hat{\bm{\alpha}})}}{\sum_{i=1}^{n}\frac{A_{i}}{e_{i}(\hat{\bm{\alpha}})}}-\frac{\sum_{i=1}^{n}\frac{(1-A_{i})Y_{i}}{1-e_{i}(\hat{\bm{\alpha}})}}{\sum_{i=1}^{n}\frac{1-A_{i}}{1-e_{i}(\hat{\bm{\alpha}})}},$ (2.4)
where $\hat{\bm{\alpha}}$ is an estimator including the solution of (2.2). Through the well-known result that the IPW estimator can be described as an M-estimator, the $\hat{\theta}_{k}\,$s are obtained as the solution to the loss functions $\ell_{Y_{1}}(\theta_{1},\hat{\bm{\alpha}})$ and $\ell_{Y_{0}}(\theta_{0},\hat{\bm{\alpha}})$, where
$\ell_{Y_{1}}(\theta_{1},\bm{\alpha})$ $=\sum_{i=1}^{n}\frac{A_{i}}{e_{i}(\bm{\alpha})}\left(Y_{i}-\theta_{1}\right)^{2},$ (2.5)
and
$\ell_{Y_{0}}(\theta_{0},\bm{\alpha})$ $=\sum_{i=1}^{n}\frac{1-A_{i}}{1-e_{i}(\bm{\alpha})}\left(Y_{i}-\theta_{0}\right)^{2}.$ (2.6)
From the above discussions, the parameters of interest, namely $\bm{\alpha}$, $\theta_{1}$, and $\theta_{0}$, can be obtained as solutions to the corresponding loss functions (2.2), (2.5), and (2.6), respectively. In the following sections, we construct a Bayesian inference procedure by using the loss functions.
3 Bayesian Estimating Procedures
In this section, we propose a Bayesian inference procedure for the interesting causal effect $\tau$ based on the general Bayesian framework[3]. Specifically, we introduce prior distributions and sampling procedures for each parameter’s posterior distribution.
To construct the posterior inference procedure for $\theta_{1}$ and $\theta_{0}$, we formulate the general posterior
$p_{G}(\theta_{1},\theta_{0},\bm{\alpha},\lambda\mid\bm{Y},\bm{X},\bm{A})=p_{G_{\theta_{1}}}(\theta_{1}\mid\bm{Y},\bm{X},\bm{A},\bm{\alpha})p_{G_{\theta_{0}}}(\theta_{0}\mid\bm{Y},\bm{X},\bm{A},\bm{\alpha})p_{G_{\bm{\alpha}}}(\bm{\alpha},\lambda\mid\bm{X},\bm{A}),$ (3.1)
where
$p_{G_{\theta_{1}}}(\theta_{1}\mid\bm{Y},\bm{X},\bm{A},\bm{\alpha})=C_{1}p(\theta_{1}\mid\bm{\alpha})\exp\{-\omega\ell_{Y_{1}}(\theta_{1},\bm{\alpha})\},$ (3.2)
$p_{G_{\theta_{0}}}(\theta_{0}\mid\bm{Y},\bm{X},\bm{A},\bm{\alpha})=C_{0}p(\theta_{0}\mid\bm{\alpha})\exp\{-\omega\ell_{Y_{0}}(\theta_{0},\bm{\alpha})\},$ (3.3)
and
$p_{G_{\bm{\alpha}}}(\bm{\alpha},\lambda\mid\bm{X},\bm{A})=C_{a}p(\bm{\alpha}\mid\lambda)p(\lambda)\exp\{-\omega\ell_{A}(\bm{\alpha})\}.$ (3.4)
Normalized constants $C_{1}$, $C_{0}$, and $C_{a}$ are independent of $\theta_{1}$, $\theta_{0}$, and $\bm{\alpha}$ and $\lambda$, respectively. The priors of $\theta_{1}$, $\theta_{0}$, $\bm{\alpha}$, and $\lambda$ are denoted by $p(\theta_{1}|\bm{\alpha})$, $p(\theta_{0}|\bm{\alpha})$, and $p(\bm{\alpha}|\lambda)$, and $p(\lambda)$, respectively. The $\omega$ denotes the learning rate, and the method to determine its value is described later. Note that assuming a uniform prior for $\theta_{1}$ and $\theta_{0}$ and a Laplace (or double exponential) prior with a location parameter of 0 and fixed scale parameter of $\lambda$ for $\bm{\alpha}$, the maximum a posteriori (MAP) estimator of the general posterior (3.1), where $\omega=1$ is equivalent to the estimator obtained by the estimation procedure through loss functions (2.2), (2.5), and (2.6) described in Section 2.
While the general posteriors (3.2) and (3.3) depend on $\bm{\alpha}$ and $\lambda$, the general posterior (3.4) is independent of $\theta_{1}$ and $\theta_{0}$. Hence, we can obtain the marginal posterior of $\theta_{1}$ and $\theta_{0}$ as follows.
  • Step.1
    Obtain posterior draws $\bm{\alpha}^{(r)}$ and $\lambda^{(r)}$ from the general posterior (3.4) for each iterations $r=1,2,\ldots,R$.
  • Step.2
    Obtain posterior draws $\theta_{k}^{(r)}$ from the general posteriors (3.2) and (3.3) of $\theta_{k}$ for $k=0,1$.
The sampling of $\bm{\alpha}$ and $\lambda$ in Step.1 can be implemented using the weighted Bayesian bootstrap[23] or probabilistic programming language Stan[5]. In the implementation, we use the Gamma prior for $\lambda$. For Step.2, $\theta_{k}$ can be drawn from the normal distribution. Suppose that the prior of $\theta_{k}$ is the Normal distribution with the mean $\mu_{k}$ and precision $\tau_{k}$, denoted as $N(\mu_{k},\tau_{k})$. The general posterior of $\theta_{k}$ is expressed as
$p(\theta_{k}\mid\bm{Y},\bm{X},\bm{A},\bm{\alpha})$ $\propto\exp\left\{-\frac{\tau_{k}}{2}(\theta_{k}-\mu_{k})^{2}\right\}\exp\left\{-\frac{\omega}{2}\sum_{i=1}^{n}s_{ki}(\theta_{k}-Y_{i})^{2}\right\}$
$\propto\exp\left\{-\frac{\tilde{\tau}_{k}}{2}(\theta_{k}-\tilde{\mu}_{k})^{2}\ \right\}$
where
$s_{ki}=\frac{2A_{i}^{k}(1-A_{i})^{1-k}}{e_{i}(\bm{\alpha})^{k}\{1-e_{i}(\bm{\alpha})\}^{1-k}},~{}~{}\tilde{\mu}_{k}=\tilde{\tau}_{k}^{-1}\left(\tau_{k}\mu_{k}+\omega\sum_{i=1}^{n}s_{ki}Y_{i}\right)~{}~{}\mbox{and}~{}~{}\tilde{\tau}_{k}=\tau_{k}+\omega\sum_{i=1}^{n}s_{ki}.$
Therefore, in Step.2, we can easily obtain the posterior draws $\theta_{k}^{(r)}$ from $N(\tilde{\mu}_{k}^{(r)},\tilde{\tau}_{k}^{(r)})$, where
$\tilde{\mu}_{k}^{(r)}=\tilde{\tau}_{k}^{(r)}\left(\tau_{k}\mu_{k}+\omega\sum_{i=1}^{n}s_{ki}^{(r)}Y_{i}\right)~{}~{}\mbox{and}~{}~{}\tilde{\tau}_{k}^{(r)}=\tau_{k}+\omega\sum_{i=1}^{n}s_{ki}^{(r)},~{}~{}s_{ki}^{(r)}=\frac{2A_{i}^{k}(1-A_{i})^{1-k}}{e_{i}(\bm{\alpha}^{(r)})^{k}\{1-e_{i}(\bm{\alpha}^{(r)})\}^{1-k}}.$
Remark 1.
About Tuning the Learning Rate $\omega$.
Since the learning rate $\omega$ affects the inaccuracy of uncertainty quantification in the generalized Bayesian posterior, a number of studies exist on its tuning[7, 10, 22, 30]. Most of these studies involve ad hoc fine-tuning in a frequentist framework and are computationally complex and expensive. In contrast to these, we consider a simple and effective learning rate selection based on the posterior covariance information criterion (PCIC)[13, 14], which can be easily computed using the posterior draws from the generalized Bayesian posterior (3.1). The PCIC is a natural generalization of the widely applicable information criterion (WAIC)[33] and has the property of being an asymptotic equivalence to the Bayesian leave-one-out cross-validation[6].
The PCIC in our generalized Bayesian procedure is computed as follows:
$\begin{split}{\rm PCIC}={}&\frac{1}{n}\sum_{i=1}^{n}\frac{1}{R}\sum_{r=1}^{R}\nu(Y_{i},\bm{X}_{i}|\theta_{1}^{(r)},\theta_{0}^{(r)},\bm{\alpha}^{(r)})\\ {}&-\left[\frac{1}{n}\sum_{i=1}^{n}\frac{1}{R}\sum_{r=1}^{R}\left\{\nu(Y_{i},\bm{X}_{i}|\theta_{1}^{(r)},\theta_{0}^{(r)},\bm{\alpha}^{(r)})s(Y_{i},\bm{X}_{i}|\theta_{1}^{(r)},\theta_{0}^{(r)},\bm{\alpha}^{(r)})\right\}\right.\\ {}&\quad\quad-\left.\frac{1}{n}\sum_{i=1}^{n}\left\{\frac{1}{R}\sum_{r=1}^{R}\nu(Y_{i},\bm{X}_{i}|\theta_{1}^{(r)},\theta_{0}^{(r)},\bm{\alpha}^{(r)})\right\}\left\{\frac{1}{R}\sum_{r=1}^{R}s(Y_{i},\bm{X}_{i}|\theta_{1}^{(r)},\theta_{0}^{(r)},\bm{\alpha}^{(r)})\right\}\right]\end{split}$
where $s(\cdot)$ is a score function that is the negative sum of equations (2.2), (2.5) and (2.6), while $\nu(\cdot)$ is an arbitrary loss function for a future observation and the parameter vector $\theta_{1}$, $\theta_{0}$ and $\bm{\alpha}$. Here, we set $\nu(\cdot)=s(\cdot)$ since this implies a learning rate $\omega$ that reduces the prediction risk with respect to the loss function (2.2), (2.5), and (2.6).
The R program available to the GitHub repository for using our proposed method outputs the PCIC calculated for several learning rates (0.2, 0.5, 1.0, and 1.5 by default), and we only need the analysis result with the learning rate that gives the minimum PCIC.
4 Simulation Experiments
In this section, we compare the performance of our proposed method with that of competing methods using simulation datasets. The setting is based on Leacy and Stuart[19], Setodji et al[28], and Orihara et al.[24]. The iteration number for all simulation examples was $10,000$.
4.1 Data-generating Mechanism
First, we describe the data-generating mechanism used in the simulations. Assuming that there were ten covariates, $\bm{X}_{i}=(X_{i1},\dots,X_{i10})$. $X_{i1}$, $X_{i3}$, $X_{i6}$, and $X_{i9}$ were generated from $Ber(0.5)$ and $X_{i7}$ and $X_{i10}$ were generated from the standard normal distribution. $X_{i2}$ and $X_{i4}$ were generated from $N(x_{i6},0.1^{2})$ and $N(x_{i9},0.1^{2})$, respectively. $X_{i5}$ and $X_{i8}$ were generated from $Ber(expit\left\{0.4(2x_{i1}-1)\right\})$ and $Ber(expit\left\{0.4(2x_{i3}-1)\right\})$, respectively. In this setting, the correlations between the following combinations of covariates were approximately:
$Corr(X_{1},X_{5})=Corr(X_{3},X_{8})=0.2,\ Corr(X_{2},X_{6})=Corr(X_{4},X_{9})=0.9,$
and the correlations of the other combinations were approximately 00.
Next, we introduce the assignment mechanism of the treatment value $A_{i}$; that is, the true propensity score $e(\bm{X}_{i})={\rm Pr}\left(A=1|\bm{X}_{i}\right)=expit\left\{h_{A}\left(\bm{X}_{i}\right)\right\}$. Here, we considered six scenarios: “Small complexity and good overlap,” “Small complexity and poor overlap,” “Moderate complexity and good overlap,” “Moderate complexity and poor overlap,” “Large complexity and good overlap,” and “Large complexity and poor overlap.” For each situation, the linear predictor $h_{A}\left(\bm{X}_{i}\right)$ of the true propensity score model was set as follows:
Small complexity and good overlap
$h_{A}\left(\bm{X}_{i}\right)=\left(X_{i1},\dots,X_{i7}\right)\bm{\beta}_{11},\ \bm{\beta}_{11}=\left(0.4,0.8,-0.25,0.6,-0.4,-0.8,-0.5,0.7\right)^{\top}.$
Small complexity and poor overlap
$h_{A}\left(\bm{X}_{i}\right)=\left(X_{i1},\dots,X_{i7}\right)\bm{\beta}_{12},\ \bm{\beta}_{12}=2.5\bm{\beta}_{11},$
Moderate complexity and good overlap
$h_{A}\left(\bm{X}_{i}\right)=\left(X_{i1},\dots,X_{i7},X_{i2}^{2},X_{i1}\times X_{i3},X_{i2}\times X_{i4},X_{i4}\times X_{i5},X_{i5}\times X_{i6}\right)\bm{\beta}_{21},$
$\bm{\beta}_{21}=\left(0.6\bm{\beta}_{11}^{\top},1,0.96,-0.3,-0.48,-0.96\right)^{\top}.$
Moderate complexity and poor overlap
$h_{A}\left(\bm{X}_{i}\right)=\left(X_{i1},\dots,X_{i7},X_{i2}^{2},X_{i1}\times X_{i3},X_{i2}\times X_{i4},X_{i4}\times X_{i5},X_{i5}\times X_{i6}\right)\bm{\beta}_{22},$
$\bm{\beta}_{22}=\left(0.4\bm{\beta}_{11}^{\top},1,1.6,-0.5,-0.8,-1.6\right)^{\top}.$
Large complexity and good overlap
$h_{A}\left(\bm{X}_{i}\right)$ $=\left(X_{i1},\dots,X_{i7},X_{i1}\times X_{i3},X_{i5}\times X_{i6},sin\left(2(X_{i2}\times X_{i4})\right),cos\left(2(X_{i4}\times X_{i5})\right),\right.$
$\hskip 14.22636pt\left.\exp\left(2(X_{i2}\times X_{i4})\right),X_{i2}\times X_{i5}\times X_{i6}\right)\bm{\beta}_{31},$
$\bm{\beta}_{31}=\left(\bm{\beta}_{11}^{\top},0.4,-0.4,0.5,0.5,-0.25,-0.5\right)^{\top}.$
Large complexity and poor overlap
$h_{A}\left(\bm{X}_{i}\right)$ $=\left(X_{i1},\dots,X_{i7},X_{i1}\times X_{i3},X_{i5}\times X_{i6},sin\left(2(X_{i2}\times X_{i4})\right),cos\left(2(X_{i4}\times X_{i5})\right),\right.$
$\hskip 14.22636pt\left.\exp\left(2(X_{i2}\times X_{i4})\right),X_{i2}\times X_{i5}\times X_{i6}\right)\bm{\beta}_{32},$
$\bm{\beta}_{32}=\left(0.5\bm{\beta}_{11}^{\top},0.8,-0.8,1,1,-0.5,-1\right)^{\top}.$
For more details, see Orihara et al.[24].
Finally, we introduce the model for the outcome $Y_{i}$. This simulation considered the binary outcome and logistic regression model ${\rm Pr}\left(Y=1|\bm{X}_{i}\right)=expit\left\{h_{Y}\left(\bm{X}_{i}\right)\right\}$ where
$h_{Y}^{\top}\left(\bm{X}_{i}\right)$ $=\left(1,A_{i},A_{i}\times X_{i2},A_{i}\times X_{i4},X_{i1},\dots,X_{i4},X_{i8},\dots,X_{i10},X_{i2}^{2},\right.$
$\hskip 14.22636pt\left.X_{i1}\times X_{i3},X_{i2}\times X_{i4},X_{i4}\times X_{i8},X_{i8}\times X_{i9}\right)\bm{\xi},$ (4.1)
[…]
Under these settings, the true causal effects (i.e., ATE) is $0.152$ for all propensity score situations (a)(f).
4.2 Estimating Methods
We considered four methods for estimating the propensity score: ordinary logistic regression, covariate balancing propensity score (CBPS[16]), regularized calibration method (RCAL[31]), and the proposed Bayesian method.
To estimate the propensity score, we consider two situations: “Confounders” and “Predictors.” For the former situation, we use only the confounders to estimate the propensity score; $\left(X_{i1},X_{i2},X_{i3},X_{i4}\right)$. Thus, we can at least adjust for confounder bias. In the latter situation, we use the predictors of the true propensity score; $\left(X_{i1},X_{i2},\dots,X_{i7}\right)$. Thus, there is a chance to estimate an unbiased propensity score compared to the “Confounders” situations. However, note that adjusting for instrumental variables that are only related to the exposure variable ($X_{i5}$, $X_{i6}$, and $X_{i7}$) can result in unstable causal effect estimates[4].
As mentioned in Section 2, the CBPS is a robust parametric propensity score estimation method that considers model misspecification. In this simulation experiment, we only considered the first-order moment aligning with previous research[28]. Note that other methods also use the same moment condition.
For RCAL method proposed in Tan[31], it is necessary to decide on the tuning parameter $\lambda$, as described in Section 2 (see (2.1)). Specifically, to determine this parameter, we consider two approaches: making a deterministic decision or using Cross-Validation (CV). In the former approach, we prepare some values of $\lambda$ in advance. In the latter approach, CV is conducted with the loss function related to the propensity score (2.2). To implement RCAL, we employ the glm.regu and ate.ipw functions from the RCAL package in R.
The proposed method, called Bayesian RCAL (BRCAL), is based on the estimating procedure that includes estimating the learning rate $\omega$ by PCIC, as explained in Section 3. The program files for implementing our proposed method is stored on the author’s GitHub page.
4.3 Performance Metrics
We evaluated the various methods based on bias, root mean squared error (RMSE), coverage probability (CP), average length of Confidence (Credible) Interval (AvL), bias and RMSE ratio (BR/RR) compared to the proposed method, and boxplot of estimated ATE from 10,000 iterations. The bias and RMSE were calculated as follows:
$bias=\bar{\hat{\tau}}-\tau_{0},\ \ \ RMSE=\sqrt{\frac{1}{10000}\sum_{k=1}^{10000}\left(\hat{\tau}_{k}-\tau_{0}\right)^{2}},$
where $\bar{\hat{\tau}}=\frac{1}{10000}\sum_{k=1}^{10000}\hat{\theta}_{k}$, $\hat{\tau}_{k}$ is the estimate of each estimator and iteration, and $\tau_{0}$ ($=0.152$) is the true value of the ATE. The CP refers to the proportion of cases where the confidence interval or credible interval includes $\tau_{0}$. The AvL was calculated as the average length of the Confidence (Credible) Interval over 10,000–time iterations. Bias and RMSE ratio (BR/RR) are calculated as the ratio compared to the bias/RMSE of our proposed method.
4.4 Simulation Results
The results are summarized in Figures 1 and A.1 and Table A.1A.6 in Appendix A. Note that for Scenario (f), results are summarized only in tables to ensure visibility of figures. In the main manuscript, we only summarize the results under “Confounders” situations.
The ordinary logistic regression exhibits small bias in scenarios (a) and (b) since the true propensity score model is correctly specified. However, in scenarios (c)(e), where the true propensity score model cannot be specified, BRCAL yields attractive results for the bias compared to other methods, especially in small samples. CBPS performs similarly to BRCAL, particularly in large samples. RCAL also performs well in some scenarios, but exhibits unstable bias results in small samples.
All methods exhibit nearly identical RMSE values across all scenarios. If anything, RCAL performs slightly better in small samples.
The 95% coverage probability (CP) of BRCAL achieves the nominal level and is higher compared to other methods, for almost all scenarios. In scenario (e), BRCAL slightly underestimates the 95% CI. In small samples, all methods except for BRCAL underestimate the 95% CI in all scenarios. Notably, RCAL estimates comparatively narrow CIs in scenarios (b) and (e). However, this tendency is observed in other methods as well.
BRCAL, our proposed method, is considered the Bayesian version of RCAL. However, BRCAL’s performance is more akin to that of CBPS, especially in large samples. In small samples, BRCAL performs well in terms of bias and 95% CP compared to CBPS and RCAL.
To summarize, our proposed method demonstrates greater, or at least equal performance compared to other methods. Notably, it yields more accurate causal effect estimates and reliable inference results in small samples.
Figure 1: Summary of estimated causal effects: Bias, root mean squared error (RMSE), and 95% coverage probability (95%CP) in 10,000 iterations were summarized under “Covariate” situation. The left panel: $n=200$; the center panel: $n=500$; The right panel: $n=1000$. RCAL (CV) was used for the CV described in Section 4.2. BRCAL (PCIC) was used for the PCIC to select the learning rate $\omega$ from candidates: $(0.2,0.5,1.0,1.5)$ (see Section 4.2).
5 Illustrative Real Data Examples
In this section, we provide an illustrative example of a real data analysis conducted on the cohort study for male civil servants in London (Whitehall dataset)[26]. Specifically, we examined the causal effects of smoking on all-cause mortality. To estimate these effects, we considered three confounders: age, systolic blood pressure, and job grade. To account for the binary exposure, we coded the original data with $A_{i}=0$ for non-smokers, and $A_{i}=1$ otherwise. The study included 17,260 subjects, of which 7,157 (41.5%) were smokers. There were also 1,670 deaths (9.7%).
The summary results of the causal effects for each method are as follows: The crude mean difference is $6.86(5.97,7.75)$ (point estimate (95% CI)), the CBPS result is $5.62(5.58,5.68)$, and the RCAL result is $5.71(4.81,6.62)$. The result of our proposed method is $5.65(4.19,7.12)$ (posterior mean (95% CI)). The point estimates are similar for CBPS, RCAL, and BRCAL. The 95% CI for BRCAL is the widest among the three methods. Given that there are 17,260 subjects, the CI is considered too wide. However, the CI is expected to provide conservative results, based on the findings from the simulation experiments.
6 Discussions and Conclusions
In this manuscript, we introduce a novel Bayesian-based IPW estimator that incorporates a covariate balancing condition. Utilizing the General Bayes framework, which is gaining interest in Bayesian analysis, our proposed method does not require consideration of the full likelihood. Accordingly, it only necessitates minimal moment conditions, as is common in causal inference contexts. Our approach demonstrates reasonable performance, especially in small sample scenarios, as shown by various simulations and real data analyses.
To the best of our knowledge, this is the first method that applies the General Bayes framework to causal inference with a covariate balancing condition. As described in Section 3, the learning rate $\omega$ is estimated by PCIC. This method proves effective, as evidenced by the 95% CI aligning well with the nominal probability. However, relying on PCIC is an ad-hoc strategy lacking theoretical justification. In future work, we aim to develop a consensus method that is theoretically grounded.
In our proposed framework, the posterior sampling scheme is conducted in two stages. First, parameters related to the posterior of the propensity score model are sampled, followed by parameters for the IPW estimator. This approach corresponds to the “two-step inference” known in the Bayesian context [20, 29], which allows us to estimate causal effects as in standard causal inference contexts [17]. The inference is justified from a Bayesian perspective [20], provided the second step is likelihood-based. Bissiri et al. [3] note that the general Bayesian procedure ensures valid Bayesian inference by meeting Coherence and Rationality [2]. Future work should explore further justification when applying the general Bayesian framework for the second step.
While simpler approaches may substitute the second step with a direct plug-in of the propensity score posterior into the IPW estimator (2.4), our method is preferable. It marginalizes over the propensity score model[20], potentially mitigating misspecification issues[18]. Furthermore, a doubly robust estimator — valid if either the propensity score or outcome model is correctly specified — has been proposed within the Bayesian context[27], offering a competitive alternative to our method.
For future work, applying our approach to sequential treatment settings is crucial for estimating the ATE in situations with variable doses. While some methods incorporating covariate balancing conditions have been proposed [15, 1], a loss function akin to that described by Tan [31] has yet to be developed. Going forward, deriving such a loss function and considering a Bayesian estimation strategy following the same approach as presented in this manuscript will be essential. In sequential treatment settings, concerns about missing data, particularly in scenarios with extended follow-up periods, arise. In Appendix B, we consider the situations where one covariate is missing in point treatment scenarios. However, these methods cannot be directly applied to sequential settings. Xu et al. [35] introduce the Bayesian Additive Regression Tree method for imputing missing data as a Bayesian approach. This method, which is potentially adaptable to our proposed and future methodologies, could enable the construction of a comprehensive Bayesian method that effectively handles missing covariate data.
Acknowledgements: All the authors accept responsibility for the content of this manuscript and approve its submission.
Declaration of interest: The authors declare they have no conflicts of interest.
Data availability statement: The data used in the real data analysis cannot be made available for general use. The simulation analysis programs are available at the following URL:
  • https://github.com/t-momozaki/BRCAL
Appendix A Simulation Results
In this section, we additionally summarize the results under “Predictors” situations. Except for the bias results, a similar tendency is observed under “Confounders” situations. Furthermore, in scenario (b), there are unexpected results in terms of RMSE and 95% CP.
BRCAL and CBPS exhibit almost the same value of bias in almost all scenarios. In contrast, RCAL exhibits large, unstable bias compared to other methods. The results of all methods become consistent in large samples.
Table A.1: Summary of estimated causal effects under the scenario (a): Bias, root mean squared error (RMSE), 95% coverage probability (95%CP), average length of CI (AvL), and bias and RMSE ratio (BR/RR) compared to the bias/RMSE of our proposed method of the IPW estimator in 10,000 iterations were summarized for each situation (in “Covariate” column) and method (in the “Method” column). RCAL: CV was used for the CV described in Section 4.2. BRCAL: PCIC was used for the PCIC to select the learning rate $\omega$ from candidates: $(0.2,0.5,1.0,1.5)$ (see Section 4.2).
Covariate Method
Bias
(×$(\times 10^{-2})$
RMSE
(×$(\times 10^{-2})$
95%CP AvL BR RR
Bias
(×$(\times 10^{-2})$
RMSE
(×$(\times 10^{-2})$
95%CP AvL BR RR
Bias
(×$(\times 10^{-2})$
RMSE
(×$(\times 10^{-2})$
95%CP AvL BR RR
Confounders Logit 0.302 5.587 0.942 0.214 0.888 1.006 0.024 3.479 0.951 0.136 0.490 1.002 0.057 2.487 0.946 0.096 0.826 1.001
CBPS 0.283 5.585 0.941 0.214 0.832 1.005 0.016 3.480 0.951 0.136 0.327 1.002 0.050 2.487 0.947 0.096 0.725 1.001
RCAL CV 0.696 5.444 0.943 0.208 2.047 0.980 0.170 3.455 0.945 0.132 3.469 0.995 0.122 2.482 0.939 0.093 1.768 0.999
$\lambda=0.01$ 0.601 5.503 0.941 0.208 0.323 3.419 0.947 0.132 0.361 2.444 0.944 0.093
$\lambda=0.05$ 0.967 5.254 0.949 0.207 0.994 3.276 0.954 0.131 1.160 2.345 0.951 0.092
$\lambda=0.10$ 0.762 5.183 0.950 0.206 0.502 3.249 0.954 0.130 0.445 2.327 0.952 0.092
BRCAL PCIC 0.340 5.555 0.961 0.256 Ref. Ref. 0.049 3.473 0.964 0.146 Ref. Ref. 0.069 2.484 0.962 0.104 Ref. Ref.
$\omega=0.2$ 0.427 5.423 1.000 0.619 0.205 3.432 1.000 0.392 0.184 2.471 1.000 0.277
$\omega=0.5$ 0.426 5.511 0.999 0.392 0.099 3.459 1.000 0.248 0.103 2.481 1.000 0.175
$\omega=1.0$ 0.350 5.548 0.988 0.278 0.064 3.470 0.988 0.175 0.079 2.484 0.987 0.124
$\omega=1.5$ 0.337 5.559 0.957 0.227 0.048 3.473 0.961 0.143 0.066 2.485 0.958 0.101
Predictors Logit 0.362 6.454 0.931 0.232 1.003 1.022 0.395 3.977 0.941 0.150 1.079 1.010 0.189 2.752 0.951 0.107 1.212 1.007
CBPS 0.290 6.372 0.927 0.230 0.803 1.010 0.311 3.953 0.938 0.148 0.850 1.004 0.191 2.740 0.949 0.106 1.224 1.002
RCAL CV 0.953 5.799 0.926 0.208 2.640 0.919 0.860 3.846 0.913 0.132 2.350 0.977 0.117 2.708 0.915 0.093 0.750 0.990
$\lambda=0.01$ 1.288 6.654 0.896 0.208 0.834 3.872 0.912 0.132 0.196 2.651 0.922 0.093
$\lambda=0.05$ 0.972 5.595 0.936 0.208 1.373 3.461 0.941 0.131 1.036 2.408 0.946 0.093
$\lambda=0.10$ 0.771 5.342 0.943 0.207 0.877 3.348 0.946 0.130 0.295 2.347 0.949 0.092
BRCAL PCIC 0.361 6.312 0.932 0.231 Ref. Ref. 0.366 3.937 0.934 0.145 Ref. Ref. 0.156 2.734 0.939 0.103 Ref. Ref.
$\omega=0.2$ 0.652 5.993 1.000 0.621 0.628 3.848 1.000 0.392 0.012 2.701 1.000 0.277
$\omega=0.5$ 0.459 6.205 0.998 0.393 0.452 3.910 0.998 0.248 0.108 2.725 0.999 0.175
$\omega=1.0$ 0.389 6.285 0.974 0.278 0.386 3.931 0.974 0.175 0.144 2.732 0.975 0.124
$\omega=1.5$ 0.357 6.314 0.929 0.227 0.362 3.938 0.932 0.143 0.157 2.734 0.937 0.101
Left side: sample size is $n=200$; Center: sample size is $n=500$; Right side: sample size is $n=1000$;
Logit: ordinary logistic regression; CBPS: Covariate balancing propensity score[16]; RCAL: Calibrating method proposed by Tan[31]; BRCAL: Proposed method.
Table A.2: Summary of estimated causal effects under the scenario (b): Bias, root mean squared error (RMSE), 95% coverage probability (95%CP), average length of CI (AvL), and bias and RMSE ratio (BR/RR) compared to the bias/RMSE of our proposed method of the IPW estimator in 10,000 iterations were summarized for each situation (in “Covariate” column) and method (in the “Method” column). RCAL: CV was used for the CV described in Section 4.2. BRCAL: PCIC was used for the PCIC to select the learning late $\omega$ from candidates: $(0.2,0.5,1.0,1.5)$ (see Section 4.2).
Covariate Method
Bias
(×$(\times 10^{-2})$
RMSE
(×$(\times 10^{-2})$
95%CP AvL BR RR
Bias
(×$(\times 10^{-2})$
RMSE
(×$(\times 10^{-2})$
95%CP AvL BR RR
Bias
(×$(\times 10^{-2})$
RMSE
(×$(\times 10^{-2})$
95%CP AvL BR RR
Confounders Logit 0.093 6.197 0.933 0.231 0.441 1.005 0.221 3.841 0.947 0.148 0.786 1.002 0.053 2.699 0.950 0.105 2.650 1.000
CBPS 0.240 6.207 0.931 0.230 1.137 1.006 0.296 3.846 0.945 0.148 1.053 1.003 0.014 2.703 0.949 0.105 0.700 1.001
RCAL CV 0.337 5.901 0.920 0.208 1.597 0.957 0.066 3.768 0.919 0.132 0.235 0.983 0.112 2.676 0.920 0.093 5.600 0.991
$\lambda=0.01$ 0.340 6.118 0.912 0.208 0.027 3.713 0.924 0.132 0.176 2.601 0.928 0.093
$\lambda=0.05$ 0.690 5.431 0.942 0.208 0.945 3.367 0.947 0.131 1.263 2.362 0.952 0.093
$\lambda=0.10$ 0.654 5.175 0.950 0.205 0.759 3.236 0.952 0.129 0.972 2.267 0.957 0.091
BRCAL PCIC 0.211 6.168 0.932 0.229 Ref. Ref. 0.281 3.835 0.941 0.146 Ref. Ref. 0.020 2.700 0.946 0.105 Ref. Ref.
$\omega=0.2$ 0.001 5.932 1.000 0.621 0.149 3.777 1.000 0.392 0.086 2.680 1.000 0.277
$\omega=0.5$ 0.129 6.096 0.999 0.393 0.246 3.818 0.999 0.248 0.040 2.694 0.999 0.175
$\omega=1.0$ 0.182 6.149 0.976 0.278 0.281 3.831 0.978 0.175 0.026 2.699 0.977 0.124
$\omega=1.5$ 0.212 6.169 0.930 0.227 0.280 3.836 0.938 0.143 0.020 2.700 0.939 0.101
Predictors Logit 0.024 11.122 0.872 0.289 0.329 1.068 0.002 7.901 0.912 0.226 0.007 1.109 0.462 6.088 0.932 0.183 2.232 1.146
CBPS 0.030 10.645 0.411 1.022 0.318 7.193 0.914 0.246 1.178 1.010 0.230 5.339 0.897 0.174 1.111 1.005
RCAL CV 1.465 6.231 0.916 0.209 20.068 0.598 1.321 5.575 0.827 0.132 4.893 0.782 1.078 4.920 5.208 0.926
$\lambda=0.01$ 5.587 20.343 0.679 23.886 3.638 12.275 0.600 0.152 1.817 7.333 0.610 0.093
$\lambda=0.05$ 1.848 9.896 0.817 0.218 1.040 4.919 0.856 0.133 0.829 3.141 0.875 0.095
$\lambda=0.10$ 1.500 5.839 0.929 0.209 1.859 3.613 0.931 0.132 2.126 2.490 0.938 0.093
BRCAL PCIC 0.073 10.414 0.835 0.345 Ref. Ref. 0.270 7.125 0.784 0.199 Ref. Ref. 0.207 5.312 0.756 0.137 Ref. Ref.
$\omega=0.2$ 0.301 10.100 0.996 0.598 0.151 6.994 0.992 0.385 0.145 5.249 0.989 0.274
$\omega=0.5$ 0.110 10.431 0.924 0.376 0.235 7.113 0.914 0.243 0.195 5.304 0.896 0.173
$\omega=1.0$ 0.025 10.536 0.800 0.266 0.285 7.154 0.785 0.172 0.213 5.322 0.757 0.122
$\omega=1.5$ 0.022 10.574 0.712 0.217 0.289 7.167 0.692 0.140 0.218 5.328 0.668 0.100
Left side: sample size is $n=200$; Center: sample size is $n=500$; Right side: sample size is $n=1000$;
Logit: ordinary logistic regression; CBPS: Covariate balancing propensity score[16]; RCAL: Calibrating method proposed by Tan[31]; BRCAL: Proposed method.
Table A.3: Summary of estimated causal effects under the scenario (c): Bias, root mean squared error (RMSE), 95% coverage probability (95%CP), average length of CI (AvL), and bias and RMSE ratio (BR/RR) compared to the bias/RMSE of our proposed method of the IPW estimator in 10,000 iterations were summarized for each situation (in “Covariate” column) and method (in the “Method” column). RCAL: CV was used for the CV described in Section 4.2. BRCAL: PCIC was used for the PCIC to select the learning late $\omega$ from candidates: $(0.2,0.5,1.0,1.5)$ (see Section 4.2).
Covariate Method
Bias
(×$(\times 10^{-2})$
RMSE
(×$(\times 10^{-2})$
95%CP AvL BR RR
Bias
(×$(\times 10^{-2})$
RMSE
(×$(\times 10^{-2})$
95%CP AvL BR RR
Bias
(×$(\times 10^{-2})$
RMSE
(×$(\times 10^{-2})$
95%CP AvL BR RR
Confounders Logit 0.269 5.566 0.939 0.211 3.405 1.001 0.263 3.471 0.944 0.134 1.283 0.999 0.268 2.439 0.949 0.095 1.117 0.998
CBPS 0.295 5.584 0.939 0.212 3.734 1.004 0.271 3.480 0.945 0.135 1.322 1.002 0.267 2.446 0.950 0.095 1.113 1.001
RCAL CV 1.220 5.400 0.945 0.209 15.443 0.971 0.193 3.429 0.944 0.132 0.941 0.987 0.029 2.414 0.948 0.093 0.121 0.988
$\lambda=0.01$ 0.468 5.452 0.943 0.209 0.393 3.406 0.946 0.132 0.402 2.393 0.949 0.093
$\lambda=0.05$ 3.007 5.272 0.949 0.208 3.022 3.331 0.950 0.131 3.215 2.371 0.948 0.093
$\lambda=0.10$ 4.738 5.265 0.950 0.207 4.663 3.365 0.947 0.131 4.875 2.437 0.940 0.092
BRCAL PCIC 0.079 5.559 0.959 0.251 Ref. Ref. 0.205 3.473 0.960 0.146 Ref. Ref. 0.240 2.444 0.965 0.105 Ref. Ref.
$\omega=0.2$ 0.927 5.436 1.000 0.619 0.219 3.441 1.000 0.392 0.002 2.432 1.000 0.277
$\omega=0.5$ 0.220 5.513 1.000 0.392 0.072 3.464 1.000 0.248 0.170 2.440 1.000 0.175
$\omega=1.0$ 0.025 5.549 0.985 0.277 0.175 3.471 0.989 0.175 0.220 2.443 0.988 0.124
$\omega=1.5$ 0.131 5.560 0.957 0.226 0.207 3.474 0.959 0.143 0.240 2.444 0.960 0.101
Predictors Logit 1.433 6.321 0.930 0.226 1.222 1.021 1.695 3.802 0.948 0.146 1.141 1.012 1.522 2.712 0.942 0.103 1.110 1.006
CBPS 1.086 6.244 0.927 0.225 0.926 1.009 1.439 3.771 0.946 0.144 0.969 1.004 1.350 2.701 0.942 0.103 0.985 1.001
RCAL CV 2.698 5.855 0.925 0.210 2.300 0.946 2.409 3.749 0.927 0.132 1.622 0.998 2.088 2.712 0.916 0.093 1.523 1.006
$\lambda=0.01$ 2.503 6.361 0.906 0.209 2.333 3.702 0.930 0.132 2.005 2.624 0.924 0.094
$\lambda=0.05$ 3.256 5.464 0.942 0.210 3.588 3.351 0.953 0.132 3.464 2.429 0.944 0.093
$\lambda=0.10$ 4.537 5.325 0.948 0.208 4.894 3.336 0.953 0.131 4.781 2.451 0.940 0.092
BRCAL PCIC 1.173 6.190 0.932 0.228 Ref. Ref. 1.485 3.756 0.947 0.144 Ref. Ref. 1.371 2.697 0.940 0.103 Ref. Ref.
$\omega=0.2$ 1.829 5.878 1.000 0.621 1.779 3.674 1.000 0.392 1.514 2.666 1.000 0.277
$\omega=0.5$ 1.386 6.083 0.999 0.393 1.568 3.730 0.999 0.248 1.399 2.688 0.999 0.175
$\omega=1.0$ 1.230 6.164 0.975 0.278 1.504 3.750 0.982 0.175 1.380 2.693 0.978 0.124
$\omega=1.5$ 1.167 6.191 0.931 0.227 1.484 3.757 0.947 0.143 1.370 2.697 0.938 0.101
Left side: sample size is $n=200$; Center: sample size is $n=500$; Right side: sample size is $n=1000$;
Logit: ordinary logistic regression; CBPS: Covariate balancing propensity score[16]; RCAL: Calibrating method proposed by Tan[31]; BRCAL: Proposed method.
Table A.4: Summary of estimated causal effects under the scenario (d): Bias, root mean squared error (RMSE), 95% coverage probability (95%CP), average length of CI (AvL), and bias and RMSE ratio (BR/RR) compared to the bias/RMSE of our proposed method of the IPW estimator in 10,000 iterations were summarized for each situation (in “Covariate” column) and method (in the “Method” column). RCAL: CV was used for the CV described in Section 4.2. BRCAL: PCIC was used for the PCIC to select the learning late $\omega$ from candidates: $(0.2,0.5,1.0,1.5)$ (see Section 4.2).
Covariate Method
Bias
(×$(\times 10^{-2})$
RMSE
(×$(\times 10^{-2})$
95%CP AvL BR RR
Bias
(×$(\times 10^{-2})$
RMSE
(×$(\times 10^{-2})$
95%CP AvL BR RR
Bias
(×$(\times 10^{-2})$
RMSE
(×$(\times 10^{-2})$
95%CP AvL BR RR
Confounders Logit 0.781 5.684 0.936 0.215 1.402 1.001 0.576 3.540 0.948 0.137 1.291 0.999 0.565 2.470 0.949 0.097 1.247 0.997
CBPS 0.716 5.711 0.937 0.216 1.285 1.006 0.493 3.553 0.949 0.138 1.105 1.002 0.472 2.479 0.949 0.098 1.042 1.000
RCAL CV 0.615 5.449 0.942 0.208 1.104 0.959 0.273 3.470 0.944 0.132 0.612 0.979 0.476 2.432 0.943 0.093 1.051 0.981
$\lambda=0.01$ 0.135 5.506 0.939 0.209 0.118 3.441 0.946 0.132 0.181 2.400 0.947 0.093
$\lambda=0.05$ 2.363 5.269 0.949 0.208 2.450 3.319 0.951 0.131 2.294 2.332 0.951 0.092
$\lambda=0.10$ 4.909 5.241 0.949 0.206 5.449 3.375 0.946 0.130 5.517 2.439 0.940 0.092
BRCAL PCIC 0.557 5.679 0.952 0.235 Ref. Ref. 0.446 3.545 0.961 0.145 Ref. Ref. 0.453 2.478 0.963 0.104 Ref. Ref.
$\omega=0.2$ 0.524 5.522 1.000 0.619 0.047 3.503 1.000 0.392 0.284 2.465 1.000 0.277
$\omega=0.5$ 0.249 5.624 1.000 0.392 0.336 3.533 0.999 0.248 0.418 2.473 0.999 0.175
$\omega=1.0$ 0.497 5.665 0.986 0.277 0.421 3.543 0.987 0.175 0.443 2.477 0.986 0.124
$\omega=1.5$ 0.576 5.679 0.952 0.227 0.447 3.546 0.959 0.143 0.453 2.478 0.958 0.101
Predictors Logit 2.977 6.551 0.927 0.233 1.119 1.019 2.961 3.989 0.943 0.150 1.085 1.010 2.796 2.803 0.944 0.107 1.068 1.004
CBPS 2.617 6.493 0.925 0.232 0.984 1.010 2.709 3.965 0.940 0.149 0.992 1.004 2.603 2.796 0.945 0.107 0.995 1.001
RCAL CV 3.792 6.057 0.921 0.210 1.426 0.942 3.770 4.038 0.901 0.132 1.381 1.023 3.634 2.900 0.890 0.093 1.389 1.039
$\lambda=0.01$ 4.470 7.089 0.879 0.209 3.752 4.012 0.901 0.132 3.353 2.783 0.904 0.094
$\lambda=0.05$ 3.510 5.532 0.943 0.210 3.288 3.427 0.944 0.132 3.200 2.436 0.945 0.093
$\lambda=0.10$ 5.066 5.278 0.952 0.207 5.194 3.365 0.944 0.130 5.460 2.464 0.937 0.092
BRCAL PCIC 2.660 6.430 0.924 0.228 Ref. Ref. 2.730 3.948 0.931 0.144 Ref. Ref. 2.617 2.792 0.932 0.102 Ref. Ref.
$\omega=0.2$ 3.060 6.054 1.000 0.621 2.890 3.857 1.000 0.392 2.705 2.761 1.000 0.277
$\omega=0.5$ 2.754 6.305 0.998 0.393 2.779 3.918 0.998 0.248 2.635 2.780 0.999 0.175
$\omega=1.0$ 2.678 6.395 0.970 0.278 2.734 3.941 0.973 0.175 2.622 2.789 0.975 0.124
$\omega=1.5$ 2.660 6.431 0.923 0.227 2.730 3.948 0.930 0.143 2.616 2.792 0.930 0.101
Left side: sample size is $n=200$; Center: sample size is $n=500$; Right side: sample size is $n=1000$;
Logit: ordinary logistic regression; CBPS: Covariate balancing propensity score[16]; RCAL: Calibrating method proposed by Tan[31]; BRCAL: Proposed method.
Table A.5: Summary of estimated causal effects under the scenario (e): Bias, root mean squared error (RMSE), 95% coverage probability (95%CP), average length of CI (AvL), and bias and RMSE ratio (BR/RR) compared to the bias/RMSE of our proposed method of the IPW estimator in 10,000 iterations were summarized for each situation (in “Covariate” column) and method (in the “Method” column). RCAL: CV was used for the CV described in Section 4.2. BRCAL: PCIC was used for the PCIC to select the learning late $\omega$ from candidates: $(0.2,0.5,1.0,1.5)$ (see Section 4.2).
Covariate Method
Bias
(×$(\times 10^{-2})$
RMSE
(×$(\times 10^{-2})$
[…]
Bias
(×$(\times 10^{-2})$
RMSE
(×$(\times 10^{-2})$
95%CP AvL BR RR
Bias
(×$(\times 10^{-2})$
RMSE
(×$(\times 10^{-2})$
95%CP AvL BR RR
Confounders Logit 2.051 6.367 0.934 0.238 1.184 1.000 1.318 3.965 0.943 0.153 1.295 0.994 1.541 2.814 0.944 0.108 1.259 0.994
CBPS 1.664 6.409 0.933 0.238 0.961 1.006 0.995 3.996 0.943 0.154 0.977 1.002 1.213 2.835 0.944 0.109 0.991 1.001
RCAL CV 2.357 6.146 0.909 0.208 1.361 0.965 1.184 4.028 0.897 0.132 1.163 1.010 1.236 2.900 0.894 0.093 1.010 1.024
$\lambda=0.01$ 2.514 6.647 0.886 0.209 1.160 4.022 0.897 0.132 1.200 2.837 0.900 0.093
$\lambda=0.05$ 2.592 5.780 0.926 0.208 1.856 3.604 0.929 0.131 2.029 2.568 0.928 0.093
$\lambda=0.10$ 3.908 5.392 0.941 0.205 3.712 3.412 0.938 0.129 4.173 2.474 0.934 0.091
BRCAL PCIC 1.732 6.369 0.928 0.229 Ref. Ref. 1.018 3.987 0.932 0.147 Ref. Ref. 1.224 2.831 0.934 0.106 Ref. Ref.
$\omega=0.2$ 2.140 6.126 1.000 0.621 1.191 3.922 1.000 0.392 1.301 2.812 1.000 0.277
$\omega=0.5$ 1.834 6.290 0.999 0.393 1.067 3.966 0.998 0.248 1.247 2.826 0.999 0.175
$\omega=1.0$ 1.766 6.350 0.970 0.278 1.032 3.982 0.972 0.175 1.227 2.829 0.971 0.124
$\omega=1.5$ 1.730 6.369 0.926 0.227 1.016 3.987 0.927 0.143 1.226 2.832 0.925 0.101
Predictors Logit 7.883 8.079 0.916 0.269 1.337 1.028 4.404 5.067 0.934 0.182 1.173 1.028 3.593 3.566 0.942 0.131 1.239 1.018
CBPS 5.841 7.933 0.926 0.291 0.990 1.010 3.731 4.952 0.921 0.177 0.994 1.004 2.896 3.513 0.933 0.129 0.998 1.003
RCAL CV 7.678 7.404 0.854 0.208 1.302 0.942 5.147 5.324 0.820 0.132 1.371 1.080 4.384 3.964 0.788 0.093 1.511 1.131
$\lambda=0.01$ 21.105 19.361 0.722 372.870 7.462 6.823 0.741 0.133 4.864 4.211 0.760 0.093
$\lambda=0.05$ 9.754 9.576 0.802 0.339 3.206 4.024 0.904 0.132 2.347 2.800 0.907 0.094
$\lambda=0.10$ 5.735 5.976 0.915 0.208 4.067 3.472 0.939 0.130 3.832 2.485 0.934 0.092
BRCAL PCIC 5.898 7.858 0.855 0.230 Ref. Ref. 3.754 4.930 0.860 0.148 Ref. Ref. 2.901 3.504 0.860 0.105 Ref. Ref.
$\omega=0.2$ 6.123 7.423 1.000 0.619 3.808 4.786 1.000 0.392 2.955 3.451 1.000 0.277
$\omega=0.5$ 5.985 7.724 0.986 0.391 3.770 4.886 0.988 0.248 2.914 3.487 0.988 0.175
$\omega=1.0$ 5.912 7.830 0.922 0.276 3.755 4.920 0.925 0.175 2.909 3.500 0.923 0.124
$\omega=1.5$ 5.894 7.863 0.850 0.225 3.753 4.932 0.851 0.143 2.900 3.505 0.850 0.101
Left side: sample size is $n=200$; Center: sample size is $n=500$; Right side: sample size is $n=1000$;
Logit: ordinary logistic regression; CBPS: Covariate balancing propensity score[16]; RCAL: Calibrating method proposed by Tan[31]; BRCAL: Proposed method;
a: For CBPS, CIs were calculated from the results where errors did not occur. The number of errors was $\frac{2}{10000}(=0.02\%)$ when “Covariate” is Predictors.
Table A.6: Summary of estimated causal effects under the scenario (f): Bias, root mean squared error (RMSE), 95% coverage probability (95%CP), average length of CI (AvL), and bias and RMSE ratio (BR/RR) compared to the bias/RMSE of our proposed method of the IPW estimator in 10,000 iterations were summarized for each situation (in “Covariate” column) and method (in the “Method” column). RCAL: CV was used for the CV described in Section 4.2. BRCAL: PCIC was used for the PCIC to select the learning late $\omega$ from candidates: $(0.2,0.5,1.0,1.5)$ (see Section 4.2).
Covariate Method
Bias
(×$(\times 10^{-2})$
RMSE
(×$(\times 10^{-2})$
95%CP AvL BR RR
Bias
(×$(\times 10^{-2})$
RMSE
(×$(\times 10^{-2})$
95%CP AvL BR RR
Bias
(×$(\times 10^{-2})$
RMSE
(×$(\times 10^{-2})$
95%CP AvL BR RR
Confounders Logit 4.437 6.726 0.925 0.246 1.181 0.973 3.686 4.109 0.939 0.158 1.279 0.970 3.677 2.954 0.940 0.112 1.316 0.974
CBPS 3.680 6.953 0.923 0.254 0.980 1.006 2.846 4.247 0.941 0.165 0.988 1.003 2.782 3.038 0.942 0.117 0.995 1.001
RCAL CV 5.432 6.756 0.875 0.208 1.446 0.978 4.008 4.455 0.858 0.132 1.391 1.052 3.619 3.310 0.841 0.093 1.295 1.091
$\lambda=0.01$ 6.961 8.604 0.815 4.349 4.128 4.665 0.841 0.132 3.539 3.279 0.845 0.093
$\lambda=0.05$ 5.203 6.638 0.883 0.217 4.180 3.950 0.900 0.131 4.027 2.832 0.898 0.093
$\lambda=0.10$ 5.983 5.829 0.917 0.206 5.767 3.615 0.925 0.130 5.800 2.647 0.915 0.092
BRCAL PCIC 3.757 6.910 0.904 0.231 Ref. Ref. 2.882 4.236 0.915 0.148 Ref. Ref. 2.795 3.034 0.917 0.107 Ref. Ref.
$\omega=0.2$ 4.150 6.661 1.000 0.620 3.078 4.172 1.000 0.391 2.902 3.014 1.000 0.277
$\omega=0.5$ 3.869 6.832 0.997 0.392 2.937 4.214 0.997 0.247 2.819 3.026 0.996 0.175
$\omega=1.0$ 3.769 6.891 0.955 0.277 2.893 4.231 0.962 0.175 2.802 3.033 0.956 0.124
$\omega=1.5$ 3.750 6.911 0.897 0.226 2.876 4.236 0.905 0.143 2.794 3.034 0.903 0.101
Predictors Logit 7.883 8.079 0.916 0.269 1.337 1.028 7.660 4.960 0.936 0.179 1.342 1.010 8.304 3.612 0.929 0.129 1.308 1.018
CBPS 5.841 7.933 0.926 0.291 0.990 1.010 5.700 4.931 0.938 0.185 0.998 1.004 6.347 3.556 0.936 0.133 1.000 1.002
RCAL CV 7.678 7.404 0.854 0.208 1.302 0.942 11.113 6.732 0.733 0.136 1.947 1.371 13.423 5.528 2.114 1.558
$\lambda=0.01$ 21.105 19.361 0.722 372.870 16.553 9.576 0.642 0.161 15.519 6.210 0.618 0.100
$\lambda=0.05$ 9.754 9.576 0.802 0.339 6.723 4.842 0.829 0.132 6.724 3.388 0.834 0.093
$\lambda=0.10$ 5.735 5.976 0.915 0.208 5.170 3.695 0.920 0.131 5.655 2.709 0.912 0.093
BRCAL PCIC 5.898 7.858 0.855 0.230 Ref. Ref. 5.709 4.911 0.860 0.145 Ref. Ref. 6.350 3.549 0.860 0.104 Ref. Ref.
$\omega=0.2$ 6.123 7.423 1.000 0.619 5.732 4.784 1.000 0.391 6.376 3.506 1.000 0.276
$\omega=0.5$ 5.985 7.724 0.986 0.391 5.707 4.872 0.989 0.247 6.359 3.535 0.985 0.175
$\omega=1.0$ 5.912 7.830 0.922 0.276 5.708 4.902 0.924 0.175 6.351 3.546 0.919 0.123
$\omega=1.5$ 5.894 7.863 0.850 0.225 5.709 4.911 0.853 0.143 6.349 3.549 0.851 0.101
Left side: sample size is $n=200$; Center: sample size is $n=500$; Right side: sample size is $n=1000$;
Logit: ordinary logistic regression; CBPS: Covariate balancing propensity score[16]; RCAL: Calibrating method proposed by Tan[31]; BRCAL: Proposed method.
Figure A.1: Summary of estimated causal effects: Bias, root mean squared error (RMSE), and 95% coverage probability (95%CP) in 10,000 iterations were summarized under “Predictor” situation. The left panel: $n=200$; the center panel: $n=500$; The right panel: $n=1000$. RCAL (CV) was used the CV described in Section 4.2. BRCAL (PCIC) was used the PCIC to select the learning late $\omega$ from candidates: $(0.2,0.5,1.0,1.5)$ (see Section 4.2).
Appendix B Proposed Method under some Covariate Missing
In this section, we address the scenario where one covariate is missing, with the missing mechanism being Missing At Random (MAR)[21]. In this context, we explore the covariate balancing-based estimating process for the missing covariate, applying the same process as discussed in the main manuscript. Here, our focus is solely on outlining the construction of the estimation process. To simplify the discussion, we consider the situation where only one covariate is potentially missing.
Let $R_{i}$ denote the missing indicator: $R_{i}=1$ means observed and $R_{i}=0$ means missing. Also, let $\bm{X}_{i}=(R_{i}X_{1i},\bm{X}_{2i})\in\mathbb{R}^{1}\times\mathbb{R}^{p-1}$ denote the (observed) confounders. Under this setting, the observed data is described by the five pairs: $(A_{i},R_{i},Y_{i},R_{i}X_{i1},\bm{X}_{2i})$. The missing mechanism is denoted as $e^{R}_{i}(\bm{X}_{2i})={\rm Pr}(R=1|\bm{X}_{2i})$. Additionally, the propensity score is redefined as $e^{A}_{i}(R_{i},\bm{X}_{i})={\rm Pr}(A=1|R_{i},\bm{X}_{i})$. Under the following two independence assumptions, the IPW estimator for the treatment group (i.e., $\theta_{1}$), denoted by
$\hat{\theta}_{1}=\frac{\sum_{i=1}^{n}\frac{R_{i}A_{i}Y_{i}}{e^{R}_{i}e^{A}_{i}}}{\sum_{i=1}^{n}\frac{R_{i}A_{i}}{e^{R}_{i}e^{A}_{i}}},$
becomes a consistent estimator:
$(Y_{1},Y_{0})\mathop{\perp\!\!\!\perp}(A,R)|\bm{X},\ \ \ R\mathop{\perp\!\!\!\perp}X_{1}|\bm{X}_{2}.$ (B.1)
The former assumption is a straightforward extension of the strongly ignorable treatment assignment, and the latter is considered the MAR assumption (i.e., the potentially missing variable is independent of the missing indicator). Thus,
$\frac{\sum_{i=1}^{n}\frac{R_{i}A_{i}Y_{i}}{e^{R}_{i}e^{A}_{i}}}{\sum_{i=1}^{n}\frac{R_{i}A_{i}}{e^{R}_{i}e^{A}_{i}}}\stackrel{{\scriptstyle P}}{{\to}}\frac{{\rm E}\left[\frac{RAY}{e^{R}e^{A}}\right]}{{\rm E}\left[\frac{RA}{e^{R}e^{A}}\right]}=\frac{{\rm E}\left[{\rm E}\left[\frac{RA}{e^{R}e^{A}}|\bm{X}\right]{\rm E}[Y_{1}|\bm{X}]\right]}{{\rm E}\left[{\rm E}\left[\frac{RA}{e^{R}e^{A}}|\bm{X}\right]\right]},$
and
${\rm E}\left[\frac{RA}{e^{R}e^{A}}|\bm{X}\right]={\rm E}\left[\frac{RA}{{\rm Pr}(R=1,A=1|\bm{X})}|\bm{X}\right]=1.$
Therefore, from (B.1), a similar moment conditions in the main manuscript can be considered. For instance, we can consider the (2.2), replaced with following loss functions:
$\ell_{R}(\bm{\gamma})$ $=\sum_{i=1}^{n}\left[R_{i}\exp\left\{-\bm{\gamma}^{\top}g_{1}(\bm{X}_{i})\right\}+(1-R_{i})\bm{\gamma}^{\top}g_{1}\right]$
$\ell_{A}(\bm{\alpha})$ $=\sum_{i=1}^{n}\left[A_{i}\exp\left\{-\bm{\alpha}^{\top}R_{i}g_{2}(\bm{X}_{i})\right\}+(1-A_{i})\bm{\alpha}^{\top}R_{i}g_{2}+(1-A_{i})\exp\left\{\bm{\alpha}^{\top}R_{i}g_{2}\right\}-A_{i}\bm{\alpha}^{\top}R_{i}g_{2}\right].$
Using the loss functions, we can construct the similar Bayesian estimation procedure described in the main manuscript.